OpenStudy (anonymous):

Given f(x)= x-5over 3, solve for f^-1(3).

OpenStudy (anonymous):

Help please!!!

OpenStudy (anonymous):

Anyone? @mathmath333

OpenStudy (mathmath333):

is it this \(\large\tt \begin{align} \color{black}{f(x)= x-\dfrac{5}{3}}\end{align}\)

OpenStudy (anonymous):

\[\frac{ x-5 }{ 3 }\]

OpenStudy (anonymous):

possible answers ate -1, 14, 18 and 21

OpenStudy (anonymous):

@mathmath333

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{f(x)= \dfrac{x-5}{3}\\~\\ \text{put f(x)=y}\\~~\\ y= \dfrac{x-5}{3}\\~\\ 3y= x-5\\~\\ 3y+5= x\\~\\ x=3y+5\\~\\ \text{sway x and y}\\~\\ y=3x+5\\~\\ \text{put} ~~f^{-1}(x)=y\\~~\\ f^{-1}(x)=3x+5\\~\\ f^{-1}(3)=3\times 3+5\\~\\ f^{-1}(3)=14\\~\\}\end{align}\)

OpenStudy (anonymous):

Thank you!